# Enhancing Agentic RL with Progressive Reward Shaping and Value-based Sampling Policy Optimization

### v1

@@ Abstract
Large Language Models (LLMs) empowered with Tool-Integrated Reasoning (TIR) can iteratively plan, call external tools, and integrate returned information to solve complex, long-horizon reasoning tasks. Agentic Reinforcement Learning (Agentic RL) optimizes such models over full tool-interaction trajectories, but two key challenges hinder effectiveness: (1) Sparse, non-instructive rewards, such as binary 0–1 verifiable signals, provide limited guidance for intermediate steps and slow convergence; (2) Gradient degradation in Group Relative Policy Optimization (GRPO), where identical rewards within a rollout group yield zero advantage, reducing sample efficiency and destabilizing training.
To address these challenges, we propose two complementary techniques: Progressive Reward Shaping (PRS) and Value-based Sampling Policy Optimization (VSPO). PRS is a curriculum-inspired reward design that introduces dense, stage-wise feedback — encouraging models to first master parseable and properly formatted tool calls, then optimize for factual correctness and answer quality. We instantiate PRS for short-form QA (with a length-aware BLEU to fairly score concise answers) and long-form QA (with LLM-as-a-Judge scoring to prevent reward hacking). VSPO is an enhanced GRPO variant that replaces low-value samples with prompts selected by a task-value metric balancing difficulty and uncertainty, and applies value-smoothing clipping to stabilize gradient updates.
Experiments on multiple short-form and long-form QA benchmarks show that PRS consistently outperforms traditional binary rewards, and VSPO achieves superior stability, faster convergence, and higher final performance compared to PPO, GRPO, CISPO, and SFT-only baselines. Together, PRS and VSPO yield LLM-based TIR agents that generalize better across domains.
@@ Introduction
Large language models (LLMs) have shown remarkable capabilities in complex reasoning and tool-augmented decision-making, enabling applications such as question answering, planning, and interactive agents (zhao2023survey; huang2024understanding). When deployed as Tool-Integrated Reasoning (TIR) agents, LLMs are equipped with external tools — e.g., search APIs or knowledge bases — and can iteratively plan, call tools, and integrate their outputs to produce final answers. This multi-step, interleaved reasoning-with-action process is particularly effective in long-horizon tasks such as multi-hop question answering or problem solving in dynamic environments.
Training high-performing LLM-based TIR agents goes beyond single-turn reinforcement learning (zhang2025landscape). Agentic reinforcement learning (Agentic RL) explicitly optimizes the LLM policy over full tool-interaction trajectories, with feedback signals derived from the correctness and quality of the intermediate and final outputs. However, in practice, Agentic RL for TIR faces two key challenges:
  • Sparse and non-instructive rewards: Widely used reinforcement learning with verifiable rewards (RLVR) (liu2025reinforcement) often employs binary 0-1 signals to indicate pass/fail based on the final answer. These sparse signals provide limited guidance for improving intermediate tool-use steps, offer no notion of partial progress, and can lead to inefficient exploration and unstable optimization.
  • Gradient degradation in GRPO: Group Relative Policy Optimization (GRPO) (guo2025deepseek) is a strong baseline for multi-sample policy improvement, but when all rollouts in a group receive identical rewards, their computed advantages become zero. This results in dampened gradient updates, low sample efficiency, and unstable training dynamics.
@@
Experiments on long-form and short-form QA benchmarks demonstrate that PRS consistently outperforms traditional binary reward schemes, and VSPO achieves more stable training and superior performance compared to PPO, GRPO, CISPO, and SFT-only baselines. Together, PRS and VSPO yield LLM-based TIR agents that converge faster, generalize better across domains.
@@
Tool-Integrated Reasoning (TIR) (lin2025understanding) augments LLMs by enabling them to interleave natural-language reasoning with external tool calls, such as code execution or search, thereby improving adaptability in complex tasks. Early TIR systems based on in-context learning or SFT (yao2022react; schick2023toolformer) achieve promising results but often generalize poorly to unseen tasks and new tool configurations. Recent advances employ Reinforcement Learning with Verifiable Rewards (RLVR) (liu2025reinforcement), allowing agents to learn robust tool-use strategies through exploration, outperforming purely imitative methods. RL frameworks such as ToolRL (qian2025toolrl), AutoTIR (wei2025autotir), and VTool-R1 (wu2025vtool) interleave symbolic computation with language reasoning in unified rollouts, balancing precise tool-mediated operations with flexible verbal inference.
@@
Agentic RAG (ravuru2024agentic) enhances traditional RAG frameworks by embedding autonomous agents that govern the retrieval-generation process, thereby enabling dynamic decision-making, iterative refinement, and adaptive orchestration. Agents leverage established patterns such as reflection (madaan2023self), planning (madaan2023self), tool use, and multi-agent collaboration (guo2024large) to manage retrieval strategies and integrate diverse information sources. Architectures range from single-agent systems to multi-agent and hierarchical designs, with some frameworks incorporating graph-based reasoning (Agent-G (leeagent)) for richer contextual integration. Compared to static RAG, Agentic RAG achieves greater flexibility, scalability, and contextual accuracy in domains such as healthcare, finance, and education (singh2024revolutionizing).
@@
Post-pretraining optimization of LLMs encompasses Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). SFT leverages high-quality expert trajectories to align model behavior through imitation, but often exhibits limited generalization capability on out-of-distribution inputs or tasks demanding novel reasoning patterns not present in the training data. In contrast, RL facilitates active exploration by LLMs, enabling them to identify and reinforce superior behaviors that enhance generalization performance. RL methods for LLMs can be categorized as critic-based or critic-free. Critic-based approaches like PPO (schulman2017proximal) and VAPO (yue2025vapo) use a value function to provide fine-grained feedback during training. Critic-free methods like GRPO (shao2024deepseekmath) and CISPO (chen2025minimax) use only sequence-level rewards, offering greater computational efficiency and scalability.
@@
Given a question $q$ with Related background information $h$ (e.g., historical dialogue records in dialogue system), the environment $\varepsilon$ provides access to a set of tools $\mathcal{T}=\{t_{0},t_{1},\dots t_{n}\}$, the language model $\mathcal{M}$ can proactively and iteratively interact with the environment by reasoning and calling specific tools in $\mathcal{T}$, obtaining the corresponding tool results from $\varepsilon$. This process repeats until $\mathcal{M}$ provides the final answer to $q$. This interaction process can be expressed as a multi-step trajectory $\mathbb{T}$.
@@
Based on Problem Formalization, the agentic RL training objective is:
$\max_{\pi_{\theta}}\mathbb{E}_{q\sim\mathcal{D},\mathbb{T}\sim\pi_{\theta}(\cdot\mid q;\mathcal{E})}\left[r_{\phi}(q,\mathbb{T})\right]\\ -\beta\mathbb{D}_{kl}\left[\pi_{\theta}(\mathbb{T}\mid q;\mathcal{E})\|\pi_{\mathrm{ref}}(\mathbb{T}\mid q;\mathcal{E})\right],$ (3)
@@
There are two well-established policy-gradient RL methods used in Agentic RL, Proximal Policy Optimization (PPO) (schulman2017proximal) and Group Relative Policy Optimization (GRPO) (guo2025deepseek).
@@
Reward shaping belongs to the category of reward design, which is a reinforcement learning (RL) technique that augments the original reward function with additional signals to accelerate and guide the agent’s learning process, while preserving the optimal policy(ibrahim2024comprehensive). In many real-world RL tasks, rewards are sparse, delayed, or noisy, making it difficult for agents to effectively explore the environment and converge to desirable behaviors. Reward shaping addresses these challenges by incorporating additional incentives or penalties that provide intermediate feedback, thereby encouraging exploration toward promising states and actions.
@@
Richer partial order. A reward function with richer partial ordering—also described as being more discriminative or dense—can distinguish between rollouts of similar quality rather than assigning identical rewards. For example, the commonly used Reinforcement Learning with Verifiable Rewards (RLVR) (liu2025reinforcement) typically employs a binary reward of 0 and 1. Such sparse binary rewards provide limited learning signals: the model only receives ”correct” or ”incorrect” feedback, without finer-grained ”better” or ”worse” signals to guide improvement. Furthermore, for GRPO, this sparsity increases the likelihood of generating zero-advantage samples (discussed further in  VSPO (Value-based Sampling Policy Gradient)).
More instructive. Inspired by curriculum Learning (soviany2022curriculum), we design the reward function to start with simple, fundamental rewards and progressively transition into more challenging criteria. This staged approach encourages the model to first learn basic behaviors before advancing to more complex reasoning patterns, facilitating more stable and efficient learning.
@@
The standard reward function of short-form QA is binary 0-1 reward based on EM (Exact Match) (jin2025search):
@@
Answer reward $R_{a}$. Unlike the binary EM reward in Eq. (7), we design a denser reward based on BLEU  (papineni2002bleu), named short-form BLEU, which is more suitable for short-form QA.
@@
$\begin{split}&PRS_{short}=\\ &\begin{cases}R_{process}+R_{format},&\text{if }R_{process}<1,\\ R_{process}+R_{format}+R_{judge},&\text{if }R_{process}\geq 1,\\ R_{process}+R_{format}+R_{judge}+R_{a},&\begin{aligned} &\text{if }R_{process}\geq 1\\ &\text{and }R_{judge}\geq 1\end{aligned}\end{cases}\end{split}$ (14)
@@
GRPO suffers from the gradient degradation problem when all rollouts within their group have the same reward. The resulting advantages are zero, leading to zero policy gradients. This has two negative impacts.
  • the number of valid samples in a batch decreases, increasing the variance of the corresponding gradient estimation, which leads to unstable training.
@@
To address this issue, we propose Value-based Sampling Policy Optimization (VSPO), an improved variant of GRPO. Unlike the dynamic sampling approach in DAPO (yu2025dapo), which may be computationally intensive. VSPO replaces zero advantage samples with selectively sampling in the current batch, maintaining computational efficiency while improving gradient quality.
VSPO
To ensure all samples in the batch produce effective gradients, VSPO first filters out the samples $x$ in the current batch with reward variance $Var\!\left(\{R(x,y_{i})\}_{i=1}^{G}\right)$ less than the threshold $\varepsilon$ (here we set $\varepsilon$ as $1\mathrm{e}{-}6$), then replaces them by resampling from the remaining samples.
Denote current batch as $B_{j}$, prompt $x_{i}=(h_{i},q_{i})\in B_{j}$ and the rollout group of $x$ is $G_{x_{i}}$. Let $\mu_{i}$ and $\sigma_{i}^{2}$ be the mean and variance of rewards within $G_{x_{i}}$. Then we have $B_{j}=B_{j}^{\sigma^{2}>=\varepsilon}\cup B_{j}^{\sigma^{2}<\varepsilon}$ and $B_{j}^{\sigma^{2}>=\varepsilon}\cap B_{j}^{\sigma^{2}<\varepsilon}=\emptyset$.
A naive approach would be uniform sampling from $B_{j}^{\sigma^{2}>=\varepsilon}$. However, we argue this can be improved. The insight of VSPO is ”prioritize samples with higher learning value”.
The learning value of task. During the training process of GRPO, there are two types of tasks are prone to zero advantages:
  • To easy task: all rollouts of the task achieve the same high reward, indicating the model $\mathcal{M}$ already proficient in solving this task.
  • To hard task: all rollouts receive the same low reward, which means model $\mathcal{M}$ lacks the necessary capability.
To measure the value of task $x$, we reinterpret the reward mean $\mu_{x}$ and varience $\sigma^{2}_{x}$ of rollouts group of $x$ in GRPO. $\sigma^{2}_{x}$ reflects the policy’s uncertainty in solving $x$. The higher the $\sigma^{2}_{x}$, the more gradient signals it contains. $\mu_{x}$ implies the average ability for $\mathcal{M}$ to solve $x$. The higher the $\mu_{x}$, the better $x$ is solved by $\mathcal{M}$.
Based on the above insights, we define the value score of $x$ as:
@@
where $R_{max}=\max_{x\in B,y_{i}\in G_{x}}R(x,y_{i})$ is the maximum reward of current batch $B$.
The term $(R_{\max}-\mu_{x})$ captures the difficulty of $x$ and represents the potential improvement gap between the best rollout and average performance, while $\sigma_{x}^{2}$ captures the policy’s uncertainty. This formulation balances both difficulty and uncertainty, ensuring that tasks with high learning potential receive appropriate sampling priority.
Value-based sampling Based on eaquation 16, we normalize the values using a temperature-controlled Softmax:
@@
where smaller $T$ increases preference for high-score samples, while a larger $T$ yields more uniform probabilities.

### v2

@@ Abstract
Large Language Models (LLMs) empowered with Tool-Integrated Reasoning (TIR) can iteratively plan, call external tools, and integrate returned information to solve complex, long-horizon reasoning tasks. Agentic Reinforcement Learning (Agentic RL) optimizes such models over full tool-interaction trajectories, but two key challenges hinder effectiveness: (1) Sparse, non-instructive rewards, such as binary 0–1 verifiable signals, provide limited guidance for intermediate steps and slow convergence; (2) Gradient degradation in Group Relative Policy Optimization (GRPO), where identical rewards within a rollout group yield zero advantage, which reducing sample efficiency.
To address these challenges, we propose two complementary techniques: Progressive Reward Shaping (PRS) and Value-based Sampling Policy Optimization (VSPO). PRS is a curriculum-inspired reward design that introduces dense, stage-wise feedback — encouraging models to first master parseable and properly formatted tool calls, then optimize for factual correctness and answer quality. We instantiate PRS for short-form QA (with a length-aware BLEU to fairly score concise answers) and long-form QA (with LLM-as-a-Judge scoring to prevent reward hacking). VSPO is an enhanced GRPO variant that replaces zero advantages samples with prompts selected by a task-value metric balancing difficulty and uncertainty, and applies value-smoothing clipping to stabilize gradient updates.
Experiments on multiple short-form and long-form QA benchmarks show that PRS consistently outperforms traditional binary rewards, and VSPO achieves superior stability, faster convergence, and higher final performance compared to SFT, PPO and GRPO baselines. Together, PRS and VSPO yield LLM-based TIR agents that generalize better across domains.
@@ Introduction
Large language models (LLMs) have shown remarkable capabilities in complex reasoning and tool-augmented decision-making, enabling applications such as question answering, planning, and interactive agents (Zhao et al. 2023; Huang et al. 2024). When deployed as Tool-Integrated Reasoning (TIR) agents, LLMs are equipped with external tools — e.g., search APIs or knowledge bases — and can iteratively plan, call tools, and integrate their outputs to produce final answers. This multi-step, interleaved reasoning-with-action process is particularly effective in long-horizon tasks such as multi-hop question answering or problem solving in dynamic environments.
Training high-performing LLM-based TIR agents goes beyond single-turn reinforcement learning (Zhang et al. 2025). Agentic reinforcement learning (Agentic RL) explicitly optimizes the LLM policy over full tool-interaction trajectories, with feedback signals derived from the correctness and quality of the intermediate and final outputs. However, in practice, Agentic RL for TIR faces two key challenges:
  • Sparse and non-instructive rewards: Widely used reinforcement learning with verifiable rewards (RLVR) (Liu et al. 2025) often employs binary 0-1 signals to indicate pass/fail based on the final answer. These sparse signals provide limited guidance for improving intermediate tool-use steps, offer no notion of partial progress, and can lead to inefficient exploration and unstable optimization.
  • Low sample efficiency in GRPO: Group Relative Policy Optimization (GRPO) (Guo et al. 2025) is a strong baseline for multi-sample policy improvement, but when all rollouts in a group receive identical rewards, their computed advantages become zero. This results in dampened gradient updates and low sample efficiency.
@@
Experiments on long-form and short-form QA benchmarks demonstrate that PRS consistently outperforms traditional binary reward schemes, and VSPO achieves more stable training and superior performance compared to SFT, PPO and GRPO baselines. Together, PRS and VSPO yield LLM-based TIR agents that converge faster, generalize better across domains.
@@
Tool-Integrated Reasoning (TIR) (Lin and Xu 2025) augments LLMs by enabling them to interleave natural-language reasoning with external tool calls, such as code execution or search, thereby improving adaptability in complex tasks. Early TIR systems based on in-context learning or SFT (Yao et al. 2022; Schick et al. 2023) achieve promising results but often generalize poorly to unseen tasks and new tool configurations. Recent advances employ Reinforcement Learning with Verifiable Rewards (RLVR) (Liu et al. 2025), allowing agents to learn robust tool-use strategies through exploration, outperforming purely imitative methods. RL frameworks such as ToolRL (Qian et al. 2025), AutoTIR (Wei et al. 2025), and VTool-R1 (Wu et al. 2025) interleave symbolic computation with language reasoning in unified rollouts, balancing precise tool-mediated operations with flexible verbal inference.
@@
Agentic RAG (Ravuru et al. 2024) enhances traditional RAG frameworks by embedding autonomous agents that govern the retrieval-generation process, thereby enabling dynamic decision-making, iterative refinement, and adaptive orchestration. Agents leverage established patterns such as reflection (Madaan et al. 2023), planning (Madaan et al. 2023), tool use, and multi-agent collaboration (Guo et al. 2024) to manage retrieval strategies and integrate diverse information sources. Architectures range from single-agent systems to multi-agent and hierarchical designs, with some frameworks incorporating graph-based reasoning (Agent-G (Lee et al. )) for richer contextual integration. Compared to static RAG, Agentic RAG achieves greater flexibility, scalability, and contextual accuracy in domains such as healthcare, finance, and education (Singh et al. 2024).
@@
Post-pretraining optimization of LLMs encompasses Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). SFT leverages high-quality expert trajectories to align model behavior through imitation, but often exhibits limited generalization capability on out-of-distribution inputs or tasks demanding novel reasoning patterns not present in the training data. In contrast, RL facilitates active exploration by LLMs, enabling them to identify and reinforce superior behaviors that enhance generalization performance. RL methods for LLMs can be categorized as critic-based or critic-free. Critic-based approaches like PPO (Schulman et al. 2017b) and VAPO (Yue et al. 2025) use a value function to provide fine-grained feedback during training. Critic-free methods like GRPO (Shao et al. 2024) and CISPO (Chen et al. 2025) use only sequence-level rewards, offering greater computational efficiency and scalability.
@@
Given a task $x=(h,q)$ where $q$ is the questioin and $h$ is the related background information (e.g., historical dialogue records in dialogue system), the environment $\varepsilon$ provides access to a set of tools $\mathcal{T}=\{t_{0},t_{1},\dots t_{n}\}$, the language model $\mathcal{M}$ can proactively and iteratively interact with the environment by reasoning and calling specific tools in $\mathcal{T}$, obtaining the corresponding tool results from $\varepsilon$. This process repeats until $\mathcal{M}$ provides the final answer to $x$. This interaction process can be expressed as a multi-step trajectory $\mathbb{T}$.
@@
Based on Section Problem Formalization, the Agentic RL training objective is:
$\max_{\pi_{\theta}}\mathbb{E}_{q\sim\mathcal{D},\mathbb{T}\sim\pi_{\theta}(\cdot\mid x;\mathcal{E})}\left[r_{\phi}(q,\mathbb{T})\right]\\ -\beta\mathbb{D}_{kl}\left[\pi_{\theta}(\mathbb{T}\mid x;\mathcal{E})\|\pi_{\mathrm{ref}}(\mathbb{T}\mid x;\mathcal{E})\right],$ (3)
@@
There are two well-established policy-gradient RL methods used in Agentic RL, Proximal Policy Optimization (PPO) (Schulman et al. 2017b) and Group Relative Policy Optimization (GRPO) (Guo et al. 2025).
@@
Reward shaping belongs to the category of reward design, which is a reinforcement learning (RL) technique that augments the original reward function with additional signals to accelerate and guide the agent’s learning process, while preserving the optimal policy(Ibrahim et al. 2024). In many real-world RL tasks, rewards are sparse, delayed, or noisy, making it difficult for agents to effectively explore the environment and converge to desirable behaviors. Reward shaping addresses these challenges by incorporating additional incentives or penalties that provide intermediate feedback, thereby encouraging exploration toward promising states and actions.
@@
Richer partial order. A reward function with richer partial ordering—also described as being more discriminative or dense—can distinguish between rollouts of similar quality rather than assigning identical rewards. For example, the commonly used Reinforcement Learning with Verifiable Rewards (RLVR) (Liu et al. 2025) typically employs a binary reward of 0 and 1. Such sparse binary rewards provide limited learning signals: the model only receives ”correct” or ”incorrect” feedback, without finer-grained ”better” or ”worse” signals to guide improvement. Furthermore, for GRPO, this sparsity increases the likelihood of generating zero-advantage samples.
More instructive. Inspired by curriculum Learning (Soviany et al. 2022), we design the reward function to start with simple, fundamental rewards and progressively transition into more challenging criteria. This staged approach encourages the model to first learn basic behaviors before advancing to more complex reasoning patterns, facilitating more stable and efficient learning.
@@
The standard reward function of short-form QA is binary 0-1 reward based on EM (Exact Match) (Jin et al. 2025):
@@
Answer reward $R_{a}$. Unlike the binary EM reward in Eq. (7), we design a denser reward based on BLEU  (Papineni et al. 2002), named short-form BLEU, which is more suitable for short-form QA.
@@
$\begin{split}&PRS_{long}=\\ &\begin{cases}R_{process}+R_{format},&\text{if }R_{process}<1,\\ R_{process}+R_{format}+R_{judge},&\text{if }R_{process}\geq 1,\\ R_{process}+R_{format}+R_{judge}+R_{a},&\begin{aligned} &\text{if }R_{process}\geq 1\\ &\text{and }R_{judge}\geq 1\end{aligned}\end{cases}\end{split}$ (14)
@@
The selection of reward weights is based on experience. $R_{process}$, $R_{judge}$ and $R_{a}$ are equal weights and $R_{format}$ is lower weight cause format not as important as the first three.
@@
GRPO suffers from the gradient degradation problem when all rollouts within their group have the same reward. The resulting advantages are zero, leading to zero policy gradients. This has a negative impacts:
@@
To address this issue, we propose Value-based Sampling Policy Optimization (VSPO), an improved variant of GRPO. Unlike the dynamic sampling approach in DAPO (Yu et al. 2025), which may be computationally intensive. VSPO replaces zero advantage samples with selectively sampling in the current batch, maintaining computational efficiency while improving sample utilization and model learning efficiency.
The introduction of VSPO
To ensure all samples in the batch produce effective gradients, VSPO first filters out the tasks $x$ in the current batch with reward variance $Var\!\left(\{R(x,y_{i})\}_{i=1}^{G}\right)$ less than the threshold $\varepsilon$ (here we set $\varepsilon$ as $1\mathrm{e}{-}6$), then replaces them by resampling from the remaining samples.
Denote current batch as $B$ and the rollout numbers is $n$, $\forall x=(h,q)\in B$, . the rollout group of $x$ is $G_{x}=\{(y_{i},r_{i})\}_{i=1}^{n}$, here $y_{i}$ denotes the rollout and $r_{i}$ is the corresponding reward. Let $\mu_{x}$ and $\sigma_{x}^{2}$ be the mean and variance of rewards within $G_{x}$. That is
$\mu_{x}=\frac{1}{n}\sum_{i=1}^{n}r_{i}$ (16)
$\sigma_{x}^{2}=\frac{1}{n}\sum_{i=1}^{n}\left(r_{i}-\mu_{x}\right)^{2}$ (17)
After filters out the tasks $x$ in the current batch with reward variance $\sigma_{x}^{2}$ less than the threshold $\varepsilon$, we have $B=B^{\sigma_{x}^{2}>=\varepsilon}\cup B^{\sigma_{x}^{2}<\varepsilon}$ and $B^{\sigma_{x}^{2}>=\varepsilon}\cap B^{\sigma_{x}^{2}<\varepsilon}=\emptyset$.
A naive approach would be uniform sampling from $B^{\sigma_{x}^{2}>=\varepsilon}$. However, we argue this can be improved. The insight of VSPO is ”prioritize samples with higher learning value tasks”.
The learning value of task. To measure the value of task $x$, we reinterpret the reward mean $\mu_{x}$ and varience $\sigma^{2}_{x}$ of rollouts group $G_{x}$ in GRPO.
  • $\mu_{x}$ implies the average ability for $\mathcal{M}$ to solve $x$. The higher the $\mu_{x}$, the better $x$ is solved by $\mathcal{M}$.
  • $\sigma_{x}^{2}$ reflects the policy’s uncertainty in solving $x$ or the diversity of exploration. The higher the $\sigma^{2}_{x}$, the more rollouts $G_{x}$ contains with relatively high and low rewards, that is to say, more gradient signals it contains.
We believe that the more difficult it is for $\mathcal{M}$ to solve and the more gradient signals the task contains have richer learning value. Based on the above insights, we define the value score of task $x$ as:
@@
where $R_{max}=\max_{x\in B,y_{i}\in G_{x}}R(x,y_{i})$ is the maximum reward of current batch $B$. $\mu_{x}$ and $\sigma^{2}_{x}$ is defined in equation 16 and 17.
The term $(R_{\max}-\mu_{x})$ captures the difficulty of $x$ and represents the potential improvement gap between the best rollout and average performance, while $\sigma_{x}^{2}$ captures the policy’s uncertainty. This formulation balances both difficulty and uncertainty, ensuring that the more difficult for $\mathcal{M}$ to solve and the more gradient signals the task contains have higher learning value.
Value-based sampling Based on eaquation 18, we normalize the values using a temperature-controlled Softmax:
@@
where smaller $T$ increases preference for high value score samples, while a larger $T$ yields more uniform probabilities.
@@
Handling of some boundary situations, set current batch is $B$:
  • If $\forall x\in B$ have $\sigma_{x}^{2}>\epsilon$, which means there is no zero advantage task, then value-based sampling will not be performed. The processing flow is exactly the same as traditional GRPO.
  • IF $\forall x\in B$ have $\sigma_{x}^{2}<=\epsilon$, which means every task of $B$ is zero advantage, then value-based sampling will not be performed, too. The processing flow is exactly the same as traditional GRPO.